OpenStudy (anonymous):

the sum of the digits of a two-digit number is 11 and the difference of the digits is 5. what is the number?

OpenStudy (anonymous):

let digit 1 =x let digit 2 =y x+y= 11 x-y = 5 solve it simultaneously.

OpenStudy (anonymous):

8+3=11 8-3=5